X+y=16
Y=-5x
How do I get the answer

Answer by rothauserc(4718)

x+y = 16
y = -5x
:
substitute -5x for y in the first equation
:
x-5x = 16
:
-4x = 16
:
*****************
x = -4
y = -5(-4) = 20
****************
:
check the answer by substituting for x and y in the first equation
:
-4 +20 = 16
:
16 = 16
:
our answer checks
: